TV advertising agencies face increasing challenges in reaching audience members because viewing TV programs via digital streaHelp with b) please. I would appreciate if you can show me the whole procedure. Thanks!
Answer #1

Ans b) 461

Explanation:

Sample proportion \hat{p} = 0.5

Critical value at 99% CI is 2.576

Margin of Error E \leq 0.06

E = z_{\alpha/2}\sqrt{\frac{\hat{p}(1-\hat{p})}{n}}

2a/2 n p(1-P) E

n = \left [ \frac{2.576\sqrt{0.5(1-0.5)}}{0.06} \right ]^2

n =460.8178\approx 461